OpenStudy (anonymous):

The one-way train fare from Chicago, IL, to Sedalia, MO, is $70. If the distance between the two points is 471 miles, what is the cost per mile of the trip?

OpenStudy (danjs):

looking for dollars/mile

OpenStudy (anonymous):

looking for dollars

OpenStudy (danjs):

\[\frac{ $70 }{ 1~trip }*\frac{ 1~trip }{ 471~mi } = \frac{ $X }{ mi }\] Dollars per trip , Trip per mile, The Trip unit cancels

OpenStudy (danjs):

Left with, dollars/ 1 mile

OpenStudy (danjs):

70 dollars / 471 miles , just divide to get # Dollars / 1 mile

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

welcome, those probs are easier taking care of units first, then deal with the #'s..good luck